A baseball player hit 61 homeruns in a season. Of the 61 homeruns 19 went to the right field and 16 went to right center field, 11 went to center field, and 11 went to left center field, and 4 went to left field. What is the probability that a randomly selected homerun was to hit the right field.
Find the probability P (E or F)if E and f are mutually exclusive. p(E)=.30 P(F)=.51

Answer by stanbon(75887)

A baseball player hit 61 homeruns in a season. Of the 61 homeruns 19 went to the right field and 16 went to right center field, 11 went to center field, and 11 went to left center field, and 4 went to left field. What is the probability that a randomly selected homerun was to hit the right field.
Ans: 19/61
-----------------------------
Find the probability P (E or F)if E and f are mutually exclusive. p(E)=.30 P(F)=.51
P(E or F) = P(E)+P(F) = 0.30+0.51 = 0.81
---------------
Cheers,
Stan H.
----------------